# Prototyping a High Purity Germanium cryogenic veto system for a bolometric detection experiment

### v1

@@ 3.3 Anti-coincidence analysis
The vetoing capability of the HPGe detectors was investigated by looking at time correlations with signals from the LWO bolometer. For the present setup, the rate of accidental coincidences is mainly dominated by the time resolution of the NTD sensor, which delivers much slower signals than the HPGe ionization detectors. The pulse onset time of an event in the LWO bolometer was defined by the intersection between the baseline and a straight line extrapolated from the point where the slope of the signal rise time is maximum (see figure 7 b)). Using such a definition, figure 9 shows the distribution of time differences between an event selected in the LWO bolometer and the closest event in time selected in any of the two HPGe detectors, after applying a 70- offset. It shows a clear peak centered at zero along with exponential tails at high corresponding to accidental coincidences. The correlation peak features a full width at half maximum of s. A time window (red dotted vertical lines) is set from -2 ms to +3 ms to select true coincidence events. Bolometer events selected in coincidence with HPGe events with a time difference falling within this time-window were thus rejected. Given the relatively high counting rates observed in the HPGe detectors, the probability for accidental veto rejection turned out to be significant. In order to estimate the true veto rejection factor, a correction factor taking into account accidental rejection was applied. This correction factor was obtained by extrapolating the fitted exponential side tails into the correlation time window (see figure 9). For a typical background run with Pb shielding, the fraction of LWO events in coincidence with at least one HPGe detector was found to be % before correction and % after correction. This result is in good agreement with the prediction of a simple toy MC that takes into account the measured rates of single and coincident events in the various detectors. Removing the Pb shielding, the fraction of rejected events by the HPGe veto detectors dropped at the 20% level because of the increase in the environmental gamma background at low energy.
@@
The energy spectrum for LWO events rejected using this time-coincidence selection criterion is shown on Figure 8 and is compared to the spectrum of single events. In order to determine the energy distribution for true coincidences, the LWO energy spectrum was corrected for by using events located outside the correlation time region, and normalized to the fraction of accidentally vetoed events. As expected, no gamma-ray line is observed for the set of rejected events, except for the case of the 511 keV gamma-ray line originating from the annihilation of pairs. A similar conclusion can be drawn for events originating from neutron captures on . Moreover, events in coincidence with the HPGe detectors account for most of the high-energy part of the single event spectrum, especially in the 6-20 MeV region. These events correspond to atmospheric muons simultaneously crossing the detectors.
@@ 4.1 Simulation description
A simulation of the experimental setup using the Geant 4 10.7 library GEANT4 is performed to validate the experimental results and disentangle the different source contributions to the selected single event energy spectra. As shown on figure 1, a simplified geometry has been implemented, i.e. the detectors along with their mechanical housings, the main elements of the dry refrigerator and the surrounding Pb shielding. Three main source contributions have been considered and separately simulated: (i) environmental gammas coming from natural radioactivity in the materials surrounding the setup, surface (ii) muons and (iii) neutrons both originating from the interactions of primary cosmic rays in the atmosphere. Environmental gammas are isotropically shot towards the setup with an energy spectrum obtained from the simulation of the radioactive decay of , and primordial radionuclides uniformly distributed in the concrete walls, ceiling and floor of a room. Atmospheric muon energy and angular distributions follow a modified Gaisser parametrization Tang_2006 . Atmospheric neutrons are isotropically shot following an experimental spectrum as measured in Gordon_2005 .
For this work, 10 primary muons and neutrons and 10 primary gammas have been generated, which corresponds to an equivalent acquisition time of 44 days and 31 days for neutrons and muons and 5 hours for environmental gammas, respectively. The so-called "Shielding" reference physics list released by the Geant 4 collaboration was used to model all particle interaction processes, except the electro-magnetic component for which the Livermore model was chosen instead of the standard one. For each simulated primary event, only the deposited energy in each of the three detectors is scored and saved for later analysis. As a first approximation, any time information, e.g. associated to the decay of activated radioisotopes or the deposition of energy into the active detectors, is disregarded. Therefore, only true detector coincidences generated over timescales possibly much longer than the acquisition run time are simulated.
@@ 4.2 Results
| Background contribution | Fluxes (/cm/s) | |
@@
| Atmospheric muons | from Tang_2006 | |
@@
| Atmospheric neutrons | from Gordon_2005 | |
Measured at IJCLab with a HPGe portable spectrometer (See text for details).
@@
This overall good data-to-MC agreement finally allows to confidently extrapolate the performances of such a HPGe cryogenic veto system when extended to a coverage. The background rejection performances of such a setup particularly apply to the Nucleus experiment, which currently develops a similar veto system. As anticipated, the gamma and atmospheric muon veto efficiency can respectively be improved up to 72.4 % and to 97.4 % when applying a 10 keV energy threshold as typically obtained for the HPGe detectors presented in this work. All sources of surface background considered, this then leads to a total rejection efficiency of 73 %. Lowering the energy threshold to 1 keV slightly improves the rejection efficiency to 74.2 % and 97.6 % for gamma-like and muon-like events respectively, leading to a total rejection efficiency of 75 %.
@@ 5 Conclusions
Finally, a simple Geant 4 simulation of the experimental setup allowed to achieve a good description of the data, separating and measuring the main contributions at play in the energy spectrum of selected and reconstructed single events for each of the three detectors. A multi step fit of the simulation to the data was used to infer the flux of environmental gamma-rays, atmospheric neutrons and atmospheric muons, and gave results very consistent with reported literature values at shallow depth laboratories. These measurements are of interest for modeling the expected particle backgrounds in shallow depth cryogenic detector experiments, for which the background environment is often not as well characterized as in deep underground laboratories. The good understanding of the particle source environment also allowed to reproduce with fairly good accuracy the rate and spectrum of the survived LWO events after applying anti-coincidence criteria with the HPGe detectors. This allowed to confidently extrapolate the performances of such a cryogenic veto system when extended to a coverage, giving promising rejection powers at the 70 % and 97 % level for gamma-like and muon-like events, respectively.

### v2

@@ 3.4 Anti-coincidence analysis
The vetoing capability of the HPGe detectors was investigated by looking at time correlations with signals from the LWO bolometer. For the present setup, the rate of accidental coincidences is mainly dominated by the time resolution of the NTD sensor, which delivers much slower signals than the HPGe ionization detectors. The time of an event in the LWO bolometer was estimated by the intersection between the baseline and a straight line extrapolated from the point where the slope of the pulse rise time is maximum (see figure 4 b)). Using such a definition, figure 9 shows the distribution of time differences between an event selected in the LWO bolometer and the closest event in time selected in any of the two HPGe detectors, after correcting for a 70- offset. As illustrated by the inset of figure 4, this offset likely sources from the LWO event time reconstruction method, which is systematically biased with respect to the true pulse onset time T0. The corrected time difference distribution shows a clear peak centered at zero along with exponential tails at high corresponding to accidental coincidences. The correlation peak features a full width at half maximum of s. A time window (red dotted vertical lines on figure 9) is set from -2 ms to +3 ms to select true coincidence events. Bolometer events selected in coincidence with HPGe events with a time difference falling within this time-window were thus rejected. Given the relatively high counting rates observed in the HPGe detectors, the probability for accidental veto rejection turned out to be significant. In order to estimate the true veto rejection factor, a correction factor taking into account accidental rejection was applied. This correction factor was obtained by extrapolating the fitted exponential side tails into the correlation time window (see figure 9). For a typical run with Pb shielding, the fraction of LWO events in coincidence with at least one HPGe detector was found to be % before correction and % after correction. This result is in good agreement with the prediction of a simple toy MC that takes into account the measured rates of single and coincident events in the various detectors. Removing the Pb shielding, the fraction of rejected events by the HPGe veto detectors dropped at the 20% level because of the increase in the environmental gamma background at low energy.
@@
The energy spectrum for LWO events rejected using this time-coincidence selection criterion is shown on Figure 8 and is compared to the spectrum of single events. In order to determine the energy distribution for true coincidences, the LWO energy spectrum was corrected for by using events located outside the correlation time region, and normalized to the fraction of accidentally vetoed events. As expected, no gamma-ray line is observed for the set of rejected events, except for the case of the 511 keV gamma-ray line originating from the annihilation of pairs. A similar conclusion can be drawn for events originating from neutron captures on . Moreover, events in coincidence with the HPGe detectors account for most of the high-energy part of the single event spectrum, especially in the 6-20 MeVee region. These events correspond to atmospheric muons simultaneously crossing the detectors.
@@ 4.1 Simulation description
A simulation of the experimental setup using the Geant 4 10.7 library [18] is performed to validate the experimental results and disentangle the different source contributions to the selected single event energy spectra. As shown on figure 1, a simplified geometry has been implemented, i.e. the detectors along with their mechanical housings, the main elements of the dry refrigerator and the surrounding Pb shielding. Three main source contributions have been considered and separately simulated: (i) environmental gammas coming from natural radioactivity in the materials surrounding the setup, surface (ii) muons and (iii) neutrons both originating from the interactions of primary cosmic rays in the atmosphere. They are expected to be the most contributing in a setup operated at shallow depth. Environmental gammas are isotropically shot towards the setup with an energy spectrum obtained from the simulation of the radioactive decay of , and primordial radionuclides uniformly distributed in the concrete walls, ceiling and floor of a room. Atmospheric muon energy and angular distributions follow a modified Gaisser parametrization [19]. Atmospheric neutrons are isotropically shot following an experimental spectrum as measured in [20].
For this work, 108 primary muons and neutrons and 109 primary gammas have been generated, which corresponds to an equivalent acquisition time of 44 days and 31 days for neutrons and muons and 5 hours for environmental gammas, respectively. The so-called "Shielding" reference physics list released by the Geant 4 collaboration was used to model all particle interaction processes, except the electro-magnetic component for which the Livermore model was chosen instead of the standard one. For each simulated primary event, only the deposited energy in each of the three detectors is scored and saved for later analysis. As a first approximation, any time information, e.g. associated to the decay of activated radioisotopes or the deposition of energy into the active detectors, is disregarded. Therefore, only true detector coincidences generated over timescales possibly much longer than the acquisition run time are simulated.
@@ 4.2 Results
| Background contribution | Fluxes (/cm2/s) | |
@@
| Atmospheric muons | from [19] | |
@@
| Atmospheric neutrons | from [20] | |
⋆ Measured at IJCLab with a HPGe portable spectrometer (See text for details).
@@
This overall good data-to-MC agreement finally allows to confidently extrapolate the performances of such a HPGe cryogenic veto system when extended to a coverage, which can e.g. be achieved using an hermetic arrangement of many HPGe crystals. The background rejection performances of such a setup particularly apply to the Nucleus experiment, which currently develops a similar veto system. As anticipated, the gamma and atmospheric muon veto efficiency can respectively be improved up to 73.3 % and to 92.4 % when applying a 10 keV energy threshold as typically obtained for the HPGe detectors presented in this work. All sources of surface background considered, this then leads to a total rejection efficiency of 75.8 %, which drops to 38% when restrained to the sub-keV energy range, which is the energy range of interest for the CEvNS detection at a reactor. This decrease in the total rejection power comes from atmospheric neutrons, which are poorly rejected by the HPGe cryogenic veto and whose contribution dominates in this energy regime.
Lowering the energy threshold to 1 keV slightly improves the total rejection efficiency to 74.8 % and 93.2 % for gamma-induced and muon-induced events respectively, leading to a total rejection efficiency of 77.9 % at all energies and to 56% at sub-keV energies. Achieving such low thresholds is more challenging, and would require for instance a careful design and integration of the electronic chain to mitigate and control unwanted sources of instrumental noise such as e.g. electromagnetic noise pickups and microphonics.
@@ 5 Conclusions
Finally, a simple Geant 4 simulation of the experimental setup allowed to achieve a good description of the data, separating and measuring the main contributions at play in the energy spectrum of selected and reconstructed single events for each of the three detectors. A multi step fit of the simulation to the data was used to infer the flux of environmental gamma-rays, atmospheric neutrons and atmospheric muons, and gave results very consistent with reported literature values at shallow depth laboratories. These measurements are of interest for modeling the expected particle backgrounds in shallow depth cryogenic detector experiments, for which the background environment is often not as well characterized as in deep underground laboratories. The good understanding of the particle source environment also allowed to reproduce with fairly good accuracy the rate and spectrum of the survived LWO events after applying anti-coincidence criteria with the HPGe detectors. This allowed to confidently extrapolate the performances of such a cryogenic veto system when extended to a coverage, giving promising rejection powers at the 73 % and 92 % level for gamma-like and muon-like events, respectively.